### v1

# Zero and Nonzero Energy Majorana Modes in an Extended Kitaev Chain

Abstract
This paper studies an extended Kitaev chain with three sublattices per unit cell. This extended version is obtained by hybridizing a modified Su-Schrieffer-Heeger model featuring trimerized unit cells with the standard Kitaev chain, resulting in a hexamer structure on the Majorana basis. Due to the interplay between the sublattice configuration and the $p$-wave superconducting pairing, a rich structure of edge modes beyond the expected Majorana zero modes is obtained. The various Majorana edge modes are further found to demonstrate considerable robustness against some generic perturbations and disorder. The presence of robust Majorana edge modes beyond their zero energy variations potentially offers a step forward in the ongoing efforts to unambiguously detect Majorana modes in experiments.
I Introduction
Topological superconductivity is a paradigmatic phase of matter that arises when superconducting pairing induces characteristics of topological phenomena [1, 2, 3]. Characterized by their ability to host robust edge modes that are their own antiparticles, i.e., the so called Majorana zero modes (MZMs), topological superconductors have attracted significant interest due to their potential application in fault-tolerant quantum computation [4]. Indeed, the MZMs, which are zero energy excitations that live at the boundaries of topological superconductors, may form a nonlocal qubit. Moreover, exchanging MZMs around one another, a process formally referred to as braiding, enables the execution of some quantum gate operations in a topologically protected manner [5]. Quantum computing with MZMs has since formed an active research area on its own [6, 7, 8, 5, 9, 10].
Motivated by theoretical proposals [11, 12, 13, 14], several experimental studies have reported some signatures of MZMs [15, 16, 17, 18, 19, 20, 21], though such signatures could also be attributed to other mechanisms [22, 23, 24, 25, 26]. At present, unambiguously confirming MZMs in a topological superconductor remains an experimental challenge. Nevertheless, theoretical studies of MZMs and variations of topological superconductors that host them are still actively conducted. For example, despite being Hermitian in nature, MZMs are also found to emerge in a range of non-Hermitian systems [27, 28, 29, 30, 30]. Moreover, periodically driven systems offer a richer playground that enables a different type of Majorana edge modes to emerge in addition to MZMs [31, 32, 33, 34, 35, 36]. It is expected that theoretically exploring these variations of topological superconductors may eventually lead to models whose MZMs could be more unequivocally detected experiments, which therefore serves as an important motivation for such studies.
In view of the above, this work presents another variation of topological superconductors that is based on the combination of the Kitaev chain [1] and an extended Su-Schrieffer-Heeger (SSH) model [37]. The Kitaev chain is a one-dimensional (1D) model of spinless fermions, proposed as a minimal platform to demonstrate the emergence of MZMs. Despite its simplicity, the Kitaev chain often serves as a basis for developing a more sophisticated and/or realistic topological superconducting model. Recently, a diverse range of experimental efforts has been concentrated on various platforms capable of simulating Kitaev chain under appropriate conditions, including, but not limited to, quantum dots [38, 39, 40, 41], electrical circuits [42], mechanical systems [43], and bosonic systems [44, 45].
The SSH model is one of the simplest toy models that supports a topological phase [37]. It describes a 1D chain with nearest-neighbor coupling that alternates in amplitudes between two different values, resulting in the dimerization of the lattice structure (two sites per unit cell). While mathematically very similar to the Kitaev chain, the edge modes in the SSH model are not Majorana fermionic in nature and are thus not as useful for quantum computing purposes. Nevertheless, its simplicity has led to various extensions of the SSH model that involve longer-range coupling [46, 47] and/or a larger unit cell [48, 49, 50, 51], which often result in richer topological features. For example, an extended SSH model with three sites per unit cell supports nonzero energy edge states that are symmetrically placed about zero energy [49, 50].
Drawing inspiration from the above studies of extended SSH models, efforts have been devoted to extending the Kitaev chain as well, leading to topological superconducting models with long-range hopping/pairing [52, 53, 54] and modulated structures such as dimer chains [55, 56]. Our work presents another extension to the Kitaev chain by modulating all system parameters in space with a period of three lattice sites. In the Majorana basis, this results in a lattice structure that exhibits six sites per unit cell. In our proposed model, the system parameters’ modulation is mathematically devised as a tensor product between the regular Kitaev chain [1] and the trimerized SSH chain we recently proposed [50]. By construction, the model exhibits features of long-range hopping and pairing while maintaining a relatively simple mathematical structure. In addition, it preserves fundamental symmetries, including chiral, particle-hole, and time-reversal symmetries, which are all important for its characteristics. This tapestry of particle-hole symmetry and superconducting pairing enriches the system with Majorana edge modes beyond the zero energy excitation that arises alone in the Kitaev chain and its existing extensions [52, 53, 54, 55, 56]. Moreover, the robustness of these Majorana edge modes is highlighted by their persistence under representative generic perturbations we consider below. Remarkably, we find that some of these perturbations actually enrich the system further in terms of its edge states, and therefore its topology.
The rest of this paper is organized as follows. In Sec. II.1, we introduce the model’s Hamiltonian and set up some notations. Sec. II.2 presents the momentum space description of the model, which includes its symmetry characterization, the identification of band touching points, and the construction of a suitable topological invariant. The real space description of the model is detailed in Sec. II.3, which includes the numerical calculation of the energy spectra at varying system parameters and the verification of the system’s various edge modes. In Sec. III, we briefly discuss an experimental proposal for realizing the extended Kitaev chain. In Sec. IV.1, we consider several representative perturbations and investigate how each perturbation affects the system. Section IV.2 further investigates the impact of disorder as well as its interplay with each of the previously considered perturbations. Finally, Sec. V concludes our work and outlines a number of potential research directions for future studies.
II Extended Kitaev chain
II.1 Hamiltonian
We study an extended Kitaev chain with three sites per unit cell whose Hamiltonian reads as follows:
$\mathcal{H}=H_{J_{1}}+H_{J_{2}}+H_{\delta}+H_{\Delta}+H_{\mu},$ (1)
$H_{J_{1}}$ $=2J_{1}\sum_{j=1}^{N}(c_{A,j}^{\dagger}c_{B,j}+c_{B,j}^{\dagger}c_{C,j}+\text{{h.c.}}),$ (2)
$H_{J_{2}}$ $=2J_{2}\sum_{j=1}^{N-1}(c_{A,j+1}^{\dagger}c_{B,j}+c_{B,j+1}^{\dagger}c_{C,j}+\text{{h.c.}}),$
$H_{\delta}$ $=2i\delta\sum_{j=1}^{N}(c_{B,j}^{\dagger}c_{A,j}^{\dagger}+c_{C,j}^{\dagger}c_{B,j}^{\dagger}-\text{{h.c.}}),$
$H_{\Delta}$ $=2i\Delta\sum_{j=1}^{N-1}(c_{A,j+1}^{\dagger}c_{B,j}^{\dagger}+c_{B,j+1}^{\dagger}c_{C,j}^{\dagger}-\text{{h.c.}}),$
$H_{\mu}$ $=2\mu\sum_{j=1}^{N}(c_{A,j}^{\dagger}c_{A,j}+c_{B,j}^{\dagger}c_{B,j}+c_{C,j}^{\dagger}c_{C,j}),$
where $c_{\xi,j}^{\dagger}$ ($c_{\xi,j}$) is the fermionic creation (annihilation) operator at sublattice $\xi=A,B,C$ of the $j^{\text{th}}$ unit cell, $N$ is the number of unit cells, $J_{1}$ is the the intracell hopping parameter, and $J_{2}$ is the intercell hopping parameter. The quantities $\delta$ and $\Delta$ are the intracell and intercell pairing amplitudes, respectively. $\mu$ is the chemical potential. $c_{\xi,j}^{\dagger}$ and $c_{\xi,j}$ could also be written as
$c_{\xi,j}^{\dagger}$ $=\frac{1}{2}\left(\gamma_{\xi_{1},j}-i\gamma_{\xi_{2},j}\right),$
$c_{\xi,j}$ $=\frac{1}{2}\left(\gamma_{\xi_{1},j}+i\gamma_{\xi_{2},j}\right),$
where $\gamma_{\xi_{1},j}$ and $\gamma_{\xi_{2},j}$ are two species of Majorana operators associated with each sublattice site $\xi$. Using these two relations, the Hamiltonian of Eq. 1 could be written in terms of Majorana operators
$\mathscr{H}=\mathscr{H}_{J_{1}}+\mathscr{H}_{J_{2}}+\mathscr{H}_{\delta}+\mathscr{H}_{\Delta}+\mathscr{H}_{\mu},$ (3)
$\mathscr{H}_{J_{1}}$ $=iJ_{1}\sum_{j=1}^{N}\left(\gamma_{A_{1},j}\gamma_{B_{2},j}+\gamma_{B_{1},j}\gamma_{A_{2},j}+\gamma_{C_{1},j}\gamma_{B_{2},j}\right$. (4)
$\quad\left.+\gamma_{B_{1},j}\gamma_{C_{2},j}\right),$
$\mathscr{H}_{J_{2}}$ $=iJ_{2}\sum_{j=1}^{N-1}\left(\gamma_{B_{1},j}\gamma_{A_{2},j+1}+\gamma_{A_{1},j+1}\gamma_{B_{2},j}+\gamma_{C_{1},j}\gamma_{B_{2},j+1}\right$.
$\quad\left.+\gamma_{B_{1},j+1}\gamma_{C_{2},j}\right),$
$\mathscr{H}_{\delta}$ $=i\delta\sum_{j=1}^{N}\left(\gamma_{B_{1},j}\gamma_{A_{1},j}-\gamma_{B_{2},j}\gamma_{A_{2},j}+\gamma_{C_{1},j}\gamma_{B_{1},j}\right$.
$\quad\left.-\gamma_{C_{2},j}\gamma_{B_{2},j}\right),$
$\mathscr{H}_{\Delta}$ $=i\Delta\sum_{j=1}^{N-1}\left(\gamma_{A_{1},j+1}\gamma_{B_{1},j}-\gamma_{A_{2},j+1}\gamma_{B_{2},j}+\gamma_{B_{1},j+1}\gamma_{C_{1},j}\right$.
$\quad\left.-\gamma_{B_{2},j+1}\gamma_{C_{2},j}\right),$
$\mathscr{H}_{\mu}$ $=i\mu\sum_{j=1}^{N}\left(\gamma_{A_{1},j}\gamma_{A_{2},j}+\gamma_{B_{1},j}\gamma_{B_{2},j}+\gamma_{C_{1},j}\gamma_{C_{2},j}\right).$
The effect of the various system parameters on the system’s Majorana constituents is schematically depicted in Fig. 1.
Figure 1: Schematic diagram of two adjacent unit cells of the proposed model. Each pair of Majorana sites is represented by a different color for the sake of clarity. (Top) shows the intracell hopping, intercell hopping, and chemical potential. (Bottom) shows the intracell and intercell pairings.
II.2 Momentum-space analysis
It is instructive to rewrite the Hamiltonian in momentum space, which enables a better understanding of the band structure and a clearer analysis of the topological properties. The momentum-space representation of Eq. (1) takes the form:
$\mathcal{H}$ $=\sum_{k}\psi_{k}^{\dagger}H(k)\psi_{k},$
$H(k)$ $=(\mu+(J_{1}+J_{2}\cos(k))S_{x}+J_{2}\sin(k)S_{y})\tau_{z}$
$\quad+((\delta-\Delta\cos(k))S_{y}+\Delta\sin(k)S_{x})\tau_{x},$ (5)
where $\psi_{k}=(c_{A,k},c_{B,k},c_{C,k},c_{A,-k}^{\dagger},c_{B,-k}^{\dagger},c_{C,-k}^{\dagger})^{T}$. $\tau_{\eta=x,y,z}$ is the well-known Pauli matrices, $S_{x}$ and $S_{y}$ are defined as
$S_{x}=\left(\begin{smallmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{smallmatrix}\right),\quad S_{y}=\left(\begin{smallmatrix}0&-i&0\\ i&0&-i\\ 0&i&0\end{smallmatrix}\right).$
The fermionic operators in $\psi_{k}$ are related to their real-space counterparts as follows:
$c_{\xi,k}$ $=\sum_{j}e^{-ikj}c_{\xi,j},\quad c_{\xi,k}^{\dagger}=\sum_{j}e^{ikj}c_{\xi,j}^{\dagger},$
$c_{\xi,-k}$ $=\sum_{j}e^{ikj}c_{\xi,j},\quad c_{\xi,-k}^{\dagger}=\sum_{j}e^{-ikj}c_{\xi,j}^{\dagger}.$
Before proceeding, it is helpful first to identify the symmetries of the system, which are key to understanding its behavior. We find that our model preserves chiral, particle-hole, and time-reversal symmetries, which are expressed as
$\mathcal{C}^{-1}H(k)\mathcal{C}$ $=-H(k),$ (6)
$\mathcal{P}^{-1}H(k)\mathcal{P}$ $=-H(-k),$
$\mathcal{T}^{-1}H(k)\mathcal{T}$ $=H(-k),$
where the operators $\mathcal{C}$, $\mathcal{P}$, and $\mathcal{T}$ are given as ($\mathcal{K}$ is the complex conjugation operator)
$\mathcal{C}$ $=\tau_{y}\otimes\mathbb{I}_{3\times 3},$ (7)
$\mathcal{P}$ $=(\tau_{x}\otimes\mathbb{I}_{3\times 3})\mathcal{K},$
$\mathcal{T}$ $=\mathcal{C}\mathcal{P},$
we emphasize that $\mathcal{C}$ is a unitary operator, whereas $\mathcal{P}$ and $\mathcal{T}$ are antiunitary as they involve the complex conjugation operator.
Being a combination of the regular Kitaev chain and a trimerized extended SSH model, it is expected that our model supports zero and non-zero energy Majorana edge modes. The emergence and disappearance of these edge modes are followed by an appropriate band gap closing that represents a topological phase transition. In particular, the zero energy Majorana modes (MZMs) are characterized by a band gap closing at zero energy. Such a topological phase transition could be identified by solving
[…]
this offers a straightforward way for identifying the MZMs. Due to the complexity of dealing with a $6\times 6$ matrix Hamiltonian (Eq. 5), we take into account the points of high symmetry, that is, at $k=0$ and $k=\pi$. This yields the following two conditions for band touching at $E=0$, i.e.,
$\mu$ $=$ $\sqrt{2\left((J_{1}+J_{2})^{2}-(\delta-\Delta)^{2}\right)},\quad\text{at }k=0,$
$\mu$ $=$ $\sqrt{2\left((J_{1}-J_{2})^{2}-(\delta+\Delta)^{2}\right)},\quad\text{at }k=\pi.$ (8)
Figure 2 shows some representative energy spectra in momentum space. When the system parameters satisfy Eq. (8), band touching points at $k=0$ and $k=\pm\pi$, respectively, are observed (see Figs. 2(a) and 2(b)). In contrast, at arbitrary parameter values, the system exhibits a fully gapped spectrum (see Fig. 2(c)).
Figure 2: Some representative energy spectra corresponding to Eq. (5) at $J_{1}=0.5$, $J_{2}=2.5$, and $\delta=\Delta=0.4$. (a) is at $\mu=4.24$. (b) is at $\mu=2.59$. (c) at is $\mu=3.5$.
To identify the topological nature of the MZMs, we now define and compute the associated winding number that exploits the chiral symmetry of the system. To this end, we first apply a unitary transformation that rotates the basis of Pauli matrices such that $\tau_{x}\rightarrow\tau_{y}$ and $\tau_{z}\rightarrow\tau_{x}$; this transforms the Hamiltonian of Eq. (5) into the off-diagonal matrix form
$\tilde{H}(k)=\begin{bmatrix}0&h\\ h^{\dagger}&0\end{bmatrix}$,
where
$h=\left(\begin{array}[]{ccc}\scriptstyle\mu&\scriptstyle J_{1}-\delta+(J_{2}+\Delta)\,e^{-ik}&\scriptstyle 0\\ \scriptstyle J_{1}+\delta+(J_{2}-\Delta)\,e^{ik}&\scriptstyle\mu&\scriptstyle J_{1}-\delta+(J_{2}+\Delta)\,e^{-ik}\\ \scriptstyle 0&\scriptstyle J_{1}+\delta+(J_{2}-\Delta)\,e^{ik}&\scriptstyle\mu\end{array}\right).$
In this basis, the winding number $\mathcal{W}$ can be calculated via the following expression:
$\mathcal{W}$ $=\frac{1}{2\pi i}\int_{-\pi}^{\pi}\text{Tr}\left(h^{-1}(k)\,\frac{\partial h(k)}{\partial k}\right)dk$
$=\begin{cases}\pm 1,&\text{MZMs are present}\\ 0,&\text{MZMs are absent}\end{cases},$ (9)
To demonstrate the validity of the above winding number, we shall analytically evaluate $\mathcal{W}$ at special parameter values of $J_{1}=\delta$ and $J_{2}=\Delta$. In this case, $\mathcal{W}$ takes the explicit form
$\mathcal{W}=\frac{1}{2\pi i}\oint\frac{1}{z-\frac{\mu^{2}}{8J_{1}J_{2}}}dz,$
which can be readily evaluated via residue theorem to yield
$\mathcal{W}=\begin{cases}\pm 1&\text{for }|\mu|<\sqrt{8|J_{1}J_{2}|}\\ 0&\text{for }|\mu|>\sqrt{8|J_{1}J_{2}|}\end{cases}.$ (10)
Note that this is consistent with the band touching conditions of Eq. (8), which now read $\mu=\sqrt{8|J_{1}J_{2}|}$ at the special parameter values under consideration. In the next section, we further confirm that the numerically calculated winding number above indeed agrees with the existence of MZMs at more general parameter values.
II.3 Real-space analysis and topological invariant
We start this section by presenting the energy spectra of the system in the real space. This approach yields direct observation of edge-localized states and their spatial profiles. In Fig. 3(a), we show the real-space energy spectrum, obtained by diagonalizing Eq. (1) under open boundary conditions (OBCs). The figure clearly demonstrates the presence of multiple nonzero energy edge modes at generic parameter values, which come in symmetric pairs at ($E$, E-E) around zero energy due to the chiral symmetry of the system. Moreover, at some range of $\mu$ values that yield $\mathcal{W}=\pm 1$ (see Fig. 3(b)), MZMs additionally emerge alongside these nonzero edge modes. Such MZMs are the main characteristics of the regular Kitaev chain in the topologically nontrivial regime [1]. On the other hand, nonzero energy edge modes are usually found in extended SSH models with at least three sites per unit cell [50].
That both types of edge modes coexist in our system is a consequence of our construction that mathematically combines the two models. It is also worth noting that, unlike the extended SSH models of Ref. [50], the nonzero edge modes in our system are also Majorana in nature like their zero energy counterparts. Therefore, such nonzero edge modes may potentially contribute to quantum computing applications or, at the very least, supplement the current experimental efforts in unambiguously detecting Majorana modes.
Figure 3: (a) and (c) are the energy spectrum as a function of $\mu$ and $\Delta$, respectively, for $N=60$ unit cells. (b) and (d) show the numerically calculated winding number associated with (a) and (c), respectively. (a) and (b) are evaluated at $J_{1}=0.5$, $J_{2}=2.5$, and $\delta=\Delta=0.4$. (c) and (d) are evaluated at $J_{1}=0.5$, $J_{2}=2.5$, $\delta=0.4$, and $\mu=2$. The black vertical lines mark the locations of the band touching points anticipated by Eq. (8). The red vertical dashed line indicates the value used in Fig. 4.
To gain additional insight into the nature of each Majorana edge mode observed in Fig. 3(a), we show its corresponding wavefunction profile in Fig. 4 at fixed parameter values. There, the wavefunction is defined as
$\left|\psi\right|^{2}(6j+3s+S)=\sum_{j=0}^{N-1}\sum_{s=0,1}\sum_{S=1,2,3}|\langle j+1,s,S|\psi\rangle|^{2},$ (11)
where $s\in\{0,1\}$, with $s=0$ for $c_{\xi,j}$ and $s=1$ for $c^{\dagger}_{\xi,j}$, and $S\in\{1,2,3\}$ corresponding to sublattices $A$, $B$, and $C$, respectively.
Intuitively, Eq. (11) is defined such that each combination of the sublattice, particle-hole, and lattice site degrees of freedom is assigned a unique integer, thereby allowing us to obtain the complete components of each edge mode. Figure 4 reveals that while all Majorana edge modes are highly localized, the nonzero energy edge modes only have two dominating peaks near a system’s edge (see Fig. 4(a,b,d,e)). By contrast, MZMs are found to have significant support on more degrees of freedom near a system’s edge. The considerably distinct features of the different types of edge modes in principle enable their separate experimental detection.
Figure 4: Wave function profiles associated with the system’s edge states for $N=60$ unit cells at $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, and $\mu=3.5$. Note that each subplot displays one left-localized edge state and one right-localized edge state.
III Potential experimental realization
Despite the ongoing challenges in uniquely identifying MZMs, numerous experimental efforts have successfully observed their expected signatures [15, 16, 17, 18, 19, 20, 21]. Recently, quantum dot arrays provide a controllable platform that is expected to host MZMs [38, 39, 40, 41]. In this case, a number of quantum dots can be arranged in a one-dimensional chain, with each site in the system formed by a quantum dot with adjustable energy and tunnel coupling. The hopping parameter between the adjacent quantum dots is implemented by a tunneling coupler, which provides a tunable tunnel barrier. Controlling the voltage applied to this barrier allows one to regulate the amplitude of electron tunneling between the quantum dots, thereby deciding the hopping strength [38]. Implementing the pairing parameter in quantum dot arrays is also achievable by connecting each dot to a conventional $s$-wave superconductor [57]. This approach allows superconducting correlations to be induced by the proximity effect. Although the parent superconductor provides $s$-wave pairing, the system effectively transforms to exhibit $p$-wave pairing when a combined effect of spin-orbit coupling and an external Zeeman field is applied to the system, which together splits the spin degeneracy [57]. The chemical potential in each quantum dot can be reliably controlled through gate voltages that are capacitively coupled to the dots, which allows local tweaking of the onsite energies [41]. Finally, a robust zero-bias conductance peak in this setup could be taken as a possible signature of the MZMs [15, 16, 17, 18, 19, 20, 21, 38, 39, 40, 41], whilst appropriate non-zero-bias peaks could further result from the presence of nonzero energy Majorana edge modes.
The extended Kitaev model presented in this paper is experimentally accessible with minimal adjustments to the current experimental geometry setups of quantum dot systems [38, 39, 40, 41] and other platforms [42, 43, 44, 45]. In particular, long-range hopping and pairing could be efficiently implemented as nearest-neighbor hopping and pairing by modifying the spatial arrangement of the lattice sites into a ladder-like structure [50].
IV Discussion
IV.1 Effects of perturbation
The presence and robustness of the edge modes in a noninteracting topological system are usually protected by specific symmetries. To uncover the role of each symmetry in our system, we now consider the effect of some representative perturbations that break and preserve some of these symmetries. To this end, we separately add each of the following terms to the momentum-space Hamiltonian of Eq. (5), i.e.,
[…]
where $\alpha_{\text{o}}$, $\beta_{\text{o}}$, $\zeta_{\text{o}}$, $\vartheta_{\text{o}}$ and $\varrho_{\text{o}}$ represent the corresponding perturbation strengths. The perturbations $\alpha(k)$, $\beta(k)$, and $\varrho(k)$ preserve all chiral, time-reversal, and particle-hole symmetries. Meanwhile, $\zeta(k)$ and $\vartheta(k)$ break chiral and time-reversal symmetries while preserving particle-hole symmetry. The symmetries preserved and broken by these perturbations are also summarized in Table 1.
Symmetry $\alpha(k)$ $\beta(k)$ $\zeta(k)$ $\vartheta(k)$ $\varrho(k)$
Chiral
Particle-hole
Time-reversal
Table 1: A summary of the symmetries preserved and broken by each perturbation operator discussed in Sec. IV.1
Physically, each of the first four perturbations is designed to introduce imbalance/asymmetry in one of the system parameters over each unit cell throughout the chain. Specifically, the perturbations $\alpha(k)$ and $\beta(k)$ introduce asymmetries in the intracell and intercell hopping parameters, $J_{1}$ and $J_{2}$, respectively. Meanwhile, $\zeta(k)$ and $\vartheta(k)$ introduce asymmetries in the intracell and intercell pairing parameters, $\delta$ and $\Delta$, respectively. The term $\varrho(k)$ modifies the chemical potential $\mu$. The corresponding real space expressions of the aforementioned perturbations are obtained as
$\alpha$ $=$ $2\alpha_{\text{o}}\sum_{j=1}^{N}(c_{A,j}c_{B,j}^{\dagger}+h.c.),$
$\beta$ $=$ $2\beta_{\text{o}}\sum_{j=1}^{N-1}(c_{A,j+1}^{\dagger}c_{B,j}+h.c.),$
$\zeta$ $=$ $2\zeta_{\text{o}}\sum_{j=1}^{N}(c_{B,j}^{\dagger}c_{A,j}^{\dagger}+h.c.),$
$\vartheta$ $=$ $2\vartheta_{\text{o}}\sum_{j=1}^{N-1}(c_{B,j}^{\dagger}c_{A,j+1}^{\dagger}+h.c.),$
$\varrho$ $=$ $2\varrho_{\text{o}}\sum_{j=1}^{N}(c_{A,j}^{\dagger}c_{A,j}-c_{B,j}^{\dagger}c_{B,j}-c_{C,j}^{\dagger}c_{C,j}).$
Finally, it should be noted that while the parameters for intracell and intercell pairing are originally multiplied by the imaginary component $i$ (see Eq. (1)), the inclusion of the real functions $\zeta$ and $\vartheta$ has the additional effect of making the effective pairing parameters complex valued.
Figures 5 and 6 summarize our findings by respectively plotting the energy spectrum in momentum space and real space in the presence of the aforementioned perturbations. In the momentum space spectrum of Fig. 5, the expected energies of the edge modes when OBC are applied are also plotted to highlight the spectral separation of the edge states from the bulk bands. Meanwhile, the real space spectrum of Fig. 6 explicitly highlights the fate of the system’s edge states as the perturbation strengths are varied. Together, both figures not only demonstrate the topological robustness of our system, but also highlight specific features that emerge as a result of the applied perturbations.
In Fig. 6(a), the system is under the effect of perturbation $\alpha$. As the perturbation strength $\alpha_{o}$ increases, the nonzero energy edge modes progressively merge into the bulk bands and eventually disappear. In contrast, despite the presence of zero energy band touching at some value of $\alpha_{o}$, the MZMs remain otherwise present throughout all $\alpha_{o}$ values under consideration. In Fig. 6(b), the system is under the effect of perturbation $\beta$. In particular, such a perturbation is found to yield energy splitting among the nonzero edge modes, such that some remain pinned at a fixed energy throughout the entire range of the applied perturbation strength, whilst some others quickly merge with the bulk bands. On the other hand, the MZMs remain pinned at zero energy until band closing occurs at $\beta_{o}\approx 1$, after which MZMs are absent.
In Figs. 6(c) and (d), the system is under the perturbations $\zeta$ and $\vartheta$, respectively, which both break the chiral symmetry. Qualitatively similar features are observed in both panels. First, MZMs remain present at small perturbation strengths before band closing occurs at zero energy and induces a topological phase transition. This result demonstrates that MZMs in our system are protected by the particle-hole rather than the chiral symmetry. Apart from the robustness of the MZMs, Figs. 6(c) and (d) also reveal the presence of additional nonzero edge modes that seemingly emerge from two of the bulk bands ($E\sim\pm 4$ in both panels). Such edge modes could be understood as remnants of the zero energy edge modes in the extended SSH component [50] of our model that are shifted to nonzero energy by the Kitaev component of the model. In the absence of perturbations, such edge modes overlap with some bulk bands, rendering them invisible and not topologically protected. In this case, the presence of either $\zeta$ or $\vartheta$ induces energy difference between the edge modes and the corresponding bulk bands.
Finally, Fig. 6(e) shows the energy spectrum of the system under the perturbation $\varrho$. Although the system in this case features a more complex edge mode structure, the origin of these edge modes could be easily tracked from the unperturbed case. In particular, the MZMs remain pinned at zero energy and are present before band closing at zero energy occurs.
Figure 5: Energy spectrum in momentum space in the presence of each of the five perturbations, $\alpha(k)$, $\beta(k)$, $\zeta(k)$, $\vartheta(k)$, and $\varrho(k)$. The parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, and $\mu=3.5$. Subplot (a) is at $\alpha_{o}$ = 0.5, subplot (b) is at $\beta_{o}$ = 0.5, subplot (c) is at $\zeta_{o}$= 1.5, subplot (d) is at $\vartheta_{o}$ = 1.5, and subplot (e) is at $\varrho_{o}=0.5$. The red dashed horizontal lines mark the energies of the expected edge states when OBC are applied (see Fig. 6).
Figure 6: Energy spectrum with respect to all five perturbation strengths for $N=60$ unit cells. The parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, and $\mu=3.5$. The red vertical dashed lines mark the parameter values used in Fig. 5.
IV.2 Effects of disorder
The disorder effect is unavoidable in real-world physical systems. From an experimental standpoint, it may originate from minor manufacturing flaws in the materials used or ambient noise that naturally arises due to experimental settings. In topological systems, it is typically expected that edge modes demonstrate some robustness against disorder, as long as the latter does not induce a topological phase transition due to some gap closing. This section aims to explicitly verify this.
Here, we theoretically simulate the presence of disorder by making each parameter in the system site-dependent such that $t_{j}\rightarrow t+\sigma_{j}$ (with $t=J_{1},J_{2},\delta,\Delta,\mu$), $j=1,2,\cdots,N$ and $\sigma_{j}$ are randomly taken from [X-X, $X$] ($X$ being the strength of the disorder). To this end, we introduce the disorder in each system’s parameter separately, which allows us to analyze the effect on each parameter individually. Fig. 7 summarizes our results, which indeed confirm the robustness of the edge modes in the system against the various disorders.
For completeness, in Fig. 8 we again consider the presence of each perturbation in Eq. (LABEL:eq:eq9), the strength of which is taken as site-dependent to investigate the simultaneous effect of generic perturbations and disorder. In all cases, the edge modes are found to be robust.
Figure 7: Energy spectrum with respect to the disorder strength for $N=60$ unit cells. The starting parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$ and $\mu=3.5$. (a) Disorder acts on $J_{1}$, (b) Disorder acts on $J_{2}$, (c) Disorder acts on $\delta$, (d) Disorder acts on $\Delta$, and (e) Disorder acts on $\mu$. Each data point is averaged over 100 disorder realizations.
Figure 8: Energy spectrum with respect to the disorder strength for $N=60$ unit cells. The starting parameter values for all five cases are $J_{1}=0.5$, $J_{2}=2.5$, $\delta=\Delta=0.4$, $\mu=3.5$, and $\alpha_{o}$ = $\beta_{o}$ = $\zeta_{o}$= $\vartheta_{o}$ = $\varrho_{o}$ = 0.5. (a) Disorder acts on $\alpha_{o}$, (b) Disorder acts on $\beta_{o}$, (c) Disorder acts on $\zeta_{o}$, (d) Disorder acts on $\vartheta_{o}$, and (e) Disorder acts on $\varrho_{o}$. Each data point is averaged over 100 disorder realizations.
V Conclusion
In conclusion, we have investigated an extended Kitaev chain that contains three sites per unit cell, resulting from the hybridization of a trimer SSH model [50] with the standard Kitaev chain [1], which leads to a total of six Majorana sites per unit cell. This hexamer lattice provides a richer edge state structure, which manifests itself compared to existing extended Kitaev models [52, 53, 54, 55, 56]. The system is found to preserve chiral, time-reversal, and particle-hole symmetries. To analyse the role of these symmetries in protecting the edge states, a set of generic perturbations has been applied, some of which preserve all symmetries, whereas the others break some of the symmetries. The edge modes’ resilience has been confirmed by this investigation. Furthermore, we considered the effect of spatial disorder on the system. In particular, we apply the disorder individually to each of the system parameters, including the previously introduced perturbations. As a result, we find that the system demonstrates the expected topological protection across all cases, maintaining robust edge modes.
Even though the Kitaev chain has been thoroughly studied and successfully implemented on a variety of physical platforms [38, 39, 40, 41, 42, 43, 44, 45], it remains a fundamental concept that inspires a considerable range of ongoing research. We expect that this extended version of the Kitaev chain will garner substantial attention in both theoretical and experimental contexts. From the experimental perspective, the inclusion of both robust MZMs and Majorana nonzero energy edge modes in our model provides richer experimental signatures. This coexistence of different edge modes can improve detection techniques by offering more diverse experimental signals, thereby contributing to progress in the experimental identification of Majorana modes. On the theoretical side, the model presented in this paper is amenable to further modification for exploration under the lens of non-Hermitian systems [58, 59, 60, 61, 62, 63], Floquet systems [64, 65, 51, 66, 67, 68, 69, 70, 70, 71, 72], or interacting systems [73, 74]. Another promising future aspect is to consider replacing the trimerized SSH model [50] with one that includes more sites per unit cell and similarly combining it with the Kitaev chain [1], which we expect to further enrich the edge modes’ profiles and potentially lead to a different topological structure. Finally, the technique employed for constructing our model is likely to pique the researchers’ curiosity in exploring hybridizations beyond the SSH and Kitaev paradigms, which would give rise to other families of superconductors featuring enlarged unit cells and/or different topological characteristics.
Acknowledgements.
R. W. B was supported by the Deanship of Research Oversight and Coordination (DROC) and the Interdisciplinary Research Center(IRC) for Intelligent Secure Systems (ISS) at King Fahd University of Petroleum & Minerals (KFUPM) through internal research Grant No. INSS2507.